# Expected Free Energy-based Planning as Variational Inference

### v2

@@ 1 Introduction
Despite its promise, practical implementations of EFE-based planning remain computationally demanding Kappen et al. (2012); Palmieri et al. (2022); van de Laar et al. (2024); Friston et al. (2021b); Paul et al. (2024). Existing methods often resort to approximations that undermine alignment with the FEP, particularly the foundational claim that all processing arises from variational free energy minimization. In this work, we address these limitations.
We show that EFE-based planning can be rigorously formulated as variational inference on a generative model augmented with preference and epistemic priors. Our central result demonstrates that minimizing a well-defined variational free energy functional naturally yields policies that integrate goal-directed behavior, information-seeking exploration, and bounded rationality. This formulation improves full theoretical alignment with the FEP and unifies active inference with the broader planning-as-inference paradigm, offering a scalable and principled framework for decision-making under uncertainty.
@@ 2 The Expected Free Energy Cost Function
Consider an agent described by a generative model [^2] For brevity, we omit commas in the notation for joint variables, e.g., $p(yx\theta u)=p(y,x,\theta,u)$.
$p(yx\theta u)=\underbrace{p(y|x\theta)}_{\begin{subarray}{c}\text{generation}\\ \text{of observations}\end{subarray}}\cdot\underbrace{p(x|u)}_{\begin{subarray}{c}\text{control}\\ \text{of states}\end{subarray}}\cdot\underbrace{p(\theta)}_{\begin{subarray}{c}\text{parameters}\\ \text{prior}\end{subarray}}\cdot\underbrace{p(u)}_{\begin{subarray}{c}\text{control}\\ \text{prior}\end{subarray}}\,,$ (1)
where $y$ holds observations, $x$ represents (latent) states, $\theta$ contains model parameters, and $u$ refers to a policy, i.e., a sequence of actions (controls). In the current context, all variables are considered part of a model rollout towards the future, so all variables are regarded as unobserved sequences.
@@
which describes the preferred future states. The planning objective is to infer a policy posterior $q(u)$ that, if executed, would efficiently guide the agent to these preferred states. [^3] Since $q(u)$ is conditioned solely on priors, namely the generative model $p(yx\theta u)$ and a preference prior $\hat{p}(x)$, it would be more appropriate to speak about an updated prior $q(u)$ rather than a posterior. For simplicity, in this paper, all distributions that result from inference are denoted by $q(\cdot)$ and termed posteriors.
@@
$G(u)=E_{q(yx\theta|u)}\bigg{[}\log\bigg{(}\underbrace{\frac{q(x|u)}{\hat{p}(x)}}_{\text{risk}}\cdot\underbrace{\underbrace{\frac{1}{q(y|x)}}_{\text{ambiguity}}\cdot\underbrace{\frac{q(\theta|x)}{q(\theta|yx)}}_{-\text{novelty}}}_{\text{epistemic costs}}\bigg{)}\bigg{]}\,.$ (3)
In (3), the assumption is that $q(yx\theta u)$ is the variational posterior with respect to the joint model $p(yx\theta u)\hat{p}(x)$. In section 4, we will refine this statement. Policies with a lower $G(u)$ value are regarded as more favorable. Active inference processes are based on the Free Energy Principle, a sophisticated theory grounded in core physics concepts that accounts for the behavior of living systems as if $G(u)$ were their planning cost function. On a more practical level, we summarize this theory by discussing the incentives behind the three components of $G(u)$.
@@
  • Ambiguity is the expected entropy $E_{q(x|u)}\big{[}H[q(y|x)]\big{]}$ of future observations $y$ under policy $u$. Minimizing EFE results in policies that seek well-predicted observations, leading to accurate state estimates.
@@
As an aside, if preferences would instead have been described by a distribution $\hat{p}(y)$ over desired future observations, it is common to define an alternative EFE as
$G^{\prime}(u)=E_{q(yx\theta|u)}\bigg{[}\log\bigg{(}\underbrace{\frac{1}{\hat{p}(y)}}_{\begin{subarray}{c}\text{pragmatic }\\ \text{costs}\end{subarray}}\cdot\underbrace{\underbrace{\frac{q(x|u)}{q(x|y)}}_{-\text{salience}}\cdot\underbrace{\frac{q(\theta|x)}{q(\theta|yx)}}_{-\text{novelty}}}_{\text{epistemic costs}}\bigg{)}\bigg{]}\,.$ (4)
Although $G(u)$ and $G^{\prime}(u)$ are distinct cost functions with different definitions of epistemic costs, they are centered around the same criteria. They can be regarded as specific instances of a broader policy cost function template given by
@@
where the denominator features a “biased” model $\bar{p}(yx\theta)$, incorporating both preference incentives $\hat{p}(\cdot)$ and variational posteriors $q(\cdot)$. Specifically, $G(u)$ in (3) uses $\bar{p}(yx\theta)=q(\theta|yx)q(y|x)\hat{p}(x)$, and $\bar{p}(yx\theta)$ in (4) factorizes as $\bar{p}(yx\theta)=q(\theta|yx)q(x|y)\hat{p}(y)$. In the following discussion, we focus on $G(u)$, noting that similar derivations and arguments apply to $G^{\prime}(u)$.
@@ 3 Related Work
In the Sophisticated Inference (SI) framework introduced by Friston et al. (2021b), a layer of recursive belief modeling is added to the EFE formulation. This enables deeper forms of planning, where agents consider not only “What will happen if I do this?” but also “What will I believe will happen if I do this?”—allowing for richer introspective evaluation of future outcomes.
@@ 4 EFE-based Planning as Variational Inference
Consider an agent with generative model
$p(yx\theta u)=p(y|x\theta)p(x|u)p(\theta)p(u)\,,$ (6)
@@
Consider the Variational Free Energy functional
@@
where the generative model in the denominator is augmented by both a preference prior $\hat{p}(\cdot)$ and epistemic priors $\tilde{p}(\cdot)$.
If the epistemic priors are chosen as
@@
$\tilde{p}(yx)$ $=\exp(D[q(\theta|yx),q(\theta|x)])\,,$ (9c)
then $F[q]$ decomposes as
@@
where $G(u)$ is the expected free energy as defined in (3). In (9), $H[\cdot]$ is the entropy functional and $D[\cdot,\cdot]$ is the Kullback-Leibler divergence (both defined in Appendix B).
@@
where $\sigma$ denotes the softmax function applied to the entropy $H[q(x|u)]$. Similar constructions apply to the priors $\tilde{p}(x)$ and $\tilde{p}(xy)$. In that case, the variational free energy $F[q]$ is shifted by a constant that does not affect the location of its minimum, and thus does not influence the outcome of the optimization.
@@ 5.1 Optimal Planning by Variational Inference
$F[q]$ $=E_{q(u)}\bigg{[}\log\frac{q(u)}{p(u)}+G(u)+\underbrace{E_{q(yx\theta|u)}\big{[}\log\frac{q(yx\theta|u)}{p(yx\theta|u)}\big{]}}_{=B(u)\text{ (complexity)}}\bigg{]}$
$=E_{q(u)}\bigg{[}\log\frac{q(u)}{\exp\big{(}-P(u)-G(u)-B(u)\big{)}}\bigg{]}\,,$ (12)
@@
Equation (13) is not new. A comparable formula for the optimal policy can be found in Equation 2.1 of Friston et al. (2021a). A main contribution of this paper is to demonstrate that the (previously established) optimal policy, given by (13), can be obtained through standard variational minimization of an appropriately defined free energy functional $F[q]$.
@@ 5.2 Interpretation of the Epistemic Priors
5.3 On the Complexity Term $B(u)$
In (13), $P(u)$ and $G(u)$ reflect past and future information about the policy posterior $q(u)$, respectively. The complexity term $B(u)$ represents the discrepancy (expressed as a KL divergence) between the (inferred) variational posterior $q(yx\theta|u)$ and the (ideal) Bayesian posterior $p(yx\theta|u)$.
Is $B(u)$ simply an unavoidable cost lacking benefits since, in contrast to $P(u)$ and $G(u)$, it does not provide information regarding effective policies? Not quite. Inference of $q$ must be executed on a specific platform in a particular context that provides access to a certain set of computational resources. For example, the resources available for tracking a specific car in traffic may differ depending on the overall complexity of the traffic situation. The presence of bounded computational resources can be regarded as a constraint on the inference process.
@@ 5.3 On the Complexity Term $B(u)$
While these inference constraints are not incorporated into the objective function $F[q]$ as seen in (8), the term $B(u)$ can be understood as a drive to minimize the unavoidable effects of these inference constraints. In other words, due to the complexity term $B(u)$ in (13), active inference agents that minimize (8) are Bayes-optimal planners for a given set of constraints. We refer the reader to Şenöz et al. (2021)for a discussion on how to express inference constraints explicitly in the objective function.
@@ Appendix A Proof of the Main Theorem
$=E_{q(u)}\bigg{[}\log\frac{q(u)}{p(u)}+\underbrace{E_{q(yx\theta|u)}\big{[}\log\frac{q(yx\theta|u)}{p(yx\theta|u)\hat{p}(x)\tilde{p}(u)\tilde{p}(x)\tilde{p}(yx)}\big{]}}_{C(u)}\bigg{]}$ (16b)
@@
In the above derivation, we still need to prove the transition for $C(u)$ from (16b) to (16c), which we address next. In the following, all expectations are with respect to $q(yx\theta|u)$ unless otherwise indicated.
Lemma 1 (Proof of equivalence $C(u)$ in (16b) and (16c)).
$C($ $u)=E\bigg{[}\log\frac{\overbrace{q(yx\theta|u)}^{\text{posterior}}}{\underbrace{p(yx\theta|u)}_{\text{predictive}}\underbrace{\hat{p}(x)}_{\text{utility}}\underbrace{\tilde{p}(u)\tilde{p}(x)\tilde{p}(yx)}_{\text{epistemic priors}}}\bigg{]}$ (17a)
@@
$=G(u)+\underbrace{E\bigg{[}\log\frac{q(yx\theta|u)}{p(yx\theta|u)}\bigg{]}}_{=B(u)}+\underbrace{E\bigg{[}\log\frac{q(y|x)q(\theta|yx)}{q(x|u)q(\theta|x)\tilde{p}(u)\tilde{p}(x)\tilde{p}(yx)}\bigg{]}}_{\text{choose epistemic priors to let this vanish}}$ (17c)
$=G(u)+B(u)+$ (17d)
@@
=$=\begin{cases}0&\text{if }\tilde{p}(u)=\exp(H[q(x|u)])\\ \text{const}&\text{if }\tilde{p}(u)=\sigma(H[q(x|u)])\end{cases}$ (18c)

### v4

@@ 1 Introduction
Despite its promise, practical implementations of EFE-based planning remain computationally demanding Kappen et al. (2012); Palmieri et al. (2022); van de Laar et al. (2024); Friston et al. (2021); Paul et al. (2024). Existing methods often resort to approximations that undermine alignment with the FEP, particularly the foundational claim that all processing arises from variational free energy minimization. In this work, we address these limitations.
We show that EFE-based planning can be rigorously formulated as variational inference on a generative model augmented with preference and epistemic priors. Our central result demonstrates that minimizing a well-defined variational free energy functional naturally yields policies that integrate goal-directed behavior, information-seeking exploration, and bounded rationality. This formulation improves full theoretical alignment with the FEP, and unifies active inference with the broader planning-as-inference paradigm, offering a scalable and principled framework for decision-making under uncertainty.
@@ 2 The Expected Free Energy Cost Function
Consider an agent described by a generative model $p(yx\theta u)$. [^2] For brevity, we omit commas in the notation for joint variables, e.g., $p(yx\theta u)=p(y,x,\theta,u)$. In this paper, we are only concerned with planning, so we will assume that the model predicts a sequence of future observations. A typical example of this model would be a rollout of a state space model, for instance
$p(yx\theta u)=p(x_{t})p(\theta)\underbrace{\prod_{k=t+1}^{T}p(y_{k}|x_{k},\theta)p(x_{k}|x_{k-1},u_{k})p(u_{k})}_{\text{rollout to the future}}$ (1)
where $t$ holds the current time step. In this model, $y$ denotes the sequence of future observations, $x$ represents the (latent) states, $\theta$ contains the model parameters, and $u$ refers to the policy, i.e., a sequence of future actions (controls). Because all of these variables are defined as part of a model rollout into the future, they are all treated as unobserved variables. Since (1) is designed to predict how the future is expected to unfold, we refer to it as the predictive model.
@@
which describes the preferred future states, sometimes referred to as target states. The planning objective is to infer a policy posterior $q(u)$ that, if executed, would efficiently guide the agent to these preferred states. [^3] Since $q(u)$ is conditioned solely on priors, namely the generative model $p(yx\theta u)$ and a preference prior $\hat{p}(x)$, it would be more appropriate to speak about an updated prior $q(u)$ rather than a posterior. For simplicity, in this paper, all distributions that result from inference are denoted by $q(\cdot)$ and termed posteriors.
@@
$G(u)=\underbrace{E_{q}\bigg[\log\frac{q(x|u)}{\hat{p}(x)}\bigg]}_{\text{risk}}+\underbrace{\underbrace{E_{q}\bigg[\log\frac{1}{q(y|x)}\bigg]}_{\text{ambiguity}}-\underbrace{E_{q}\bigg[\log\frac{q(\theta|yx)}{q(\theta|x)}}_{\text{novelty}}\bigg]}_{\text{epistemic costs}}\,$, (3)
where the expectations are with respect to $q=q(yx\theta|u)$. In (3), the assumption is that $q(yx\theta u)$ is the variational posterior with respect to the joint model $p(yx\theta u)\hat{p}(x)$. In section 4, we will refine this statement. Policies with a lower $G(u)$ value are regarded as more favorable, i.e., a priori more likely to be selected. Active inference processes are based on the Free Energy Principle, a sophisticated theory grounded in core physics concepts that accounts for the behavior of living systems as if $G(u)$ were their planning cost function. The most contemporary reference is Friston et al. (2023). On a more practical level, we summarize this theory by discussing the incentives behind the three components of $G(u)$.
@@
  • Ambiguity is the expected entropy $E_{q(x|u)}\big[H[q(y|x)]\big]$ of future observations $y$ under policy $u$. Minimizing EFE results in policies that seek well-predicted (i.e., unambiguous) observations, leading to accurate state estimates.
@@
As an aside, if preferences were instead described by a distribution $\hat{p}(y)$ over desired future observations, it is common to define an alternative EFE as
$G^{\prime}(u)=\underbrace{E_{q}\bigg[\log\frac{1}{\hat{p}(y)}\bigg]}_{\begin{subarray}{c}\text{pragmatic}\\ \text{costs}\end{subarray}}\underbrace{-\underbrace{E_{q}\bigg[\log\frac{q(x|y)}{q(x|u)}\bigg]}_{\text{salience}}-\underbrace{E_{q}\bigg[\log\frac{q(\theta|yx)}{q(\theta|x)}}_{\text{novelty}}\bigg]}_{\text{epistemic costs}}\,$, (4)
where, as before, the expectations are with respect to $q=q(yx\theta|u)$. Although $G(u)$ and $G^{\prime}(u)$ are distinct cost functions with different definitions of epistemic costs, they are centered around the same criteria. They can be regarded as specific instances of a broader policy cost function template given by
@@
[…]
@@ 3 Related Work
In the Sophisticated Inference (SI) framework introduced by Friston et al. (2021), a layer of recursive belief modeling is added to the EFE formulation. This enables deeper forms of planning, where agents consider not only “What will happen if I do this?” but also “What will I believe will happen if I do this?”—allowing for richer introspective evaluation of future outcomes.
@@ 4 EFE-based Planning as Variational Inference
Consider an agent with generative (predictive) model
$p(yx\theta u)\,$, (6)
@@
Let the Variational Free Energy functional $F[q]$ be defined as
@@
[…]
Let the epistemic priors be defined as
@@
$\tilde{p}(yx)$ $=\exp(D[q(\theta|yx),q(\theta|x)])\,.$ (9c)
Then, $F[q]$ decomposes as
@@
where $G(u)$ is the expected free energy as defined in (3). In (9), $H[q]=E_{q}[-\log q]$ is the entropy functional, and $D[q,p]=E_{q}[\log q-\log p]$ is the Kullback-Leibler divergence (see Appendix B).
@@
where $\sigma$ denotes the softmax function applied to the entropy $H[q(x|u)]$. Similar constructions apply to the priors $\tilde{p}(x)$ and $\tilde{p}(xy)$. In that case, the variational free energy $F[q]$ is shifted by a constant that does not affect the location of its minimum, and thus does not influence the outcome of the optimization process.
@@ 5.1 Optimal Planning by Variational Inference
$F[q]$ $=E_{q(u)}\bigg[\log\frac{q(u)}{p(u)}+G(u)+\underbrace{E_{q(yx\theta|u)}\big[\log\frac{q(yx\theta|u)}{p(yx\theta|u)}\big]}_{=C(u)\text{ (complexity)}}\bigg]$
$=E_{q(u)}\bigg[\log\frac{q(u)}{\exp\big(-P(u)-G(u)-C(u)\big)}\bigg]\,$, (12)
@@
Equation (13) is not new. A comparable formula for the optimal policy can be found in Equation 2.1 of Friston et al. (2021). A main contribution of this paper is to demonstrate that the (previously established) optimal policy, given by (13), can be obtained through standard variational minimization of an appropriately defined free energy functional $F[q]$.
@@ 5.2 Interpretation of the Epistemic Priors
5.3 On the Complexity Term $C(u)$
In (13), $P(u)$ and $G(u)$ reflect past and future information about the policy posterior $q(u)$, respectively. The complexity term $C(u)$ represents the discrepancy (expressed as a KL divergence) between the (inferred) variational posterior $q(yx\theta|u)$ and the (ideal) Bayesian posterior $p(yx\theta|u)$.
Is $C(u)$ simply an unavoidable cost lacking benefits since, in contrast to $P(u)$ and $G(u)$, it does not provide information regarding effective policies? Not quite. Inference of $q$ must be executed on a specific platform in a particular context that provides access to a certain set of computational resources. For example, the resources available for tracking a specific car in traffic may differ depending on the overall complexity of the traffic situation. The presence of bounded computational resources can be regarded as a constraint on the inference process.
@@ 5.3 On the Complexity Term $C(u)$
While these inference constraints are not incorporated into the objective function $F[q]$ as seen in (8), the term $C(u)$ can be understood as a drive to minimize the unavoidable effects of these inference constraints. In other words, due to the complexity term $C(u)$ in (13), active inference agents that minimize (8) are Bayes-optimal planners for a given set of constraints. We refer the reader to Şenöz et al. (2021)for a discussion on how to express inference constraints explicitly in the objective function.
@@ Appendix A Proof of the Main Theorem
$=E_{q(u)}\bigg[\log\frac{q(u)}{p(u)}+\underbrace{E_{q(yx\theta|u)}\big[\log\frac{q(yx\theta|u)}{p(yx\theta|u)\hat{p}(x)\tilde{p}(u)\tilde{p}(x)\tilde{p}(yx)}\big]}_{B(u)}\bigg]$ (16b)
@@
In the above derivation, we still need to prove the transition for $B(u)$ from (16b) to (16c), which we address next. In the following, all expectations are with respect to $q(yx\theta|u)$ unless otherwise indicated.
Lemma 1 (Proof of equivalence $B(u)$ in (16b) and (16c)).
$B($ $u)=E\bigg[\log\frac{\overbrace{q(yx\theta|u)}^{\text{posterior}}}{\underbrace{p(yx\theta|u)}_{\text{predictive}}\underbrace{\hat{p}(x)}_{\text{utility}}\underbrace{\tilde{p}(u)\tilde{p}(x)\tilde{p}(yx)}_{\text{epistemic priors}}}\bigg]$ (17a)
@@
$=G(u)+\underbrace{E\bigg[\log\frac{q(yx\theta|u)}{p(yx\theta|u)}\bigg]}_{=C(u)}+\underbrace{E\bigg[\log\frac{q(y|x)q(\theta|yx)}{q(x|u)q(\theta|x)\tilde{p}(u)\tilde{p}(x)\tilde{p}(yx)}\bigg]}_{\text{choose epistemic priors to let this vanish}}$ (17c)
$=G(u)+C(u)+$ (17d)
@@
$=\begin{cases}0&\text{if }\tilde{p}(u)=\exp(H[q(x|u)])\\ \text{const}&\text{if }\tilde{p}(u)=\sigma(H[q(x|u)])\end{cases}$ (18c)